\begin{thebibliography}{71}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Nazir et~al.(2020)Nazir, Rao, Wu, and Sun]{nazir2020issues}
Ambreen Nazir, Yuan Rao, Lianwei Wu, and Ling Sun.
\newblock Issues and challenges of aspect-based sentiment analysis: A comprehensive survey.
\newblock \emph{IEEE Transactions on Affective Computing}, 13\penalty0 (2):\penalty0 845--863, 2020.

\bibitem[Fei et~al.(2022{\natexlab{a}})Fei, Wu, Li, Li, Li, Qin, Zhang, Zhang, and Chua]{FeiLasuieNIPS22}
Hao Fei, Shengqiong Wu, Jingye Li, Bobo Li, Fei Li, Libo Qin, Meishan Zhang, Min Zhang, and Tat-Seng Chua.
\newblock Lasuie: Unifying information extraction with latent adaptive structure-aware generative language model.
\newblock In \emph{Proceedings of the Advances in Neural Information Processing Systems, NeurIPS 2022}, pages 15460--15475, 2022{\natexlab{a}}.

\bibitem[Kiritchenko et~al.()Kiritchenko, Zhu, Cherry, and Mohammad]{kiritchenkodetecting}
S~Kiritchenko, X~Zhu, C~Cherry, and S~Mohammad.
\newblock Detecting aspects and sentiment in customer reviews.
\newblock In \emph{8th International Workshop on Semantic Evaluation (SemEval)}, pages 437--442.

\bibitem[Dong et~al.(2014)Dong, Wei, Tan, Tang, Zhou, and Xu]{dong2014adaptive}
Li~Dong, Furu Wei, Chuanqi Tan, Duyu Tang, Ming Zhou, and Ke~Xu.
\newblock Adaptive recursive neural network for target-dependent twitter sentiment classification.
\newblock In \emph{Proceedings of the 52nd annual meeting of the association for computational linguistics (volume 2: Short papers)}, pages 49--54, 2014.

\bibitem[Zhang et~al.(2018)Zhang, Qi, and Manning]{zhang2018graph}
Yuhao Zhang, Peng Qi, and Christopher~D Manning.
\newblock Graph convolution over pruned dependency trees improves relation extraction.
\newblock \emph{EMNLP}, 2018.

\bibitem[Fei et~al.(2022{\natexlab{b}})Fei, Li, Wu, Li, Ji, and Li]{FeiDiaREIJCAI22}
Hao Fei, Jingye Li, Shengqiong Wu, Chenliang Li, Donghong Ji, and Fei Li.
\newblock Global inference with explicit syntactic and discourse structures for dialogue-level relation extraction.
\newblock In \emph{Proceedings of the Thirty-First International Joint Conference on Artificial Intelligence, {IJCAI}}, pages 4082--4088, 2022{\natexlab{b}}.

\bibitem[Wu et~al.(2021{\natexlab{a}})Wu, Fei, Ren, Ji, and Li]{Wu0RJL21}
Shengqiong Wu, Hao Fei, Yafeng Ren, Donghong Ji, and Jingye Li.
\newblock Learn from syntax: Improving pair-wise aspect and opinion terms extraction with rich syntactic knowledge.
\newblock In \emph{Proceedings of the Thirtieth International Joint Conference on Artificial Intelligence}, pages 3957--3963, 2021{\natexlab{a}}.

\bibitem[Xiang et~al.(2022)Xiang, Zhang, Li, Fei, and Ji]{xiang2022semantic}
Chunli Xiang, Junchi Zhang, Fei Li, Hao Fei, and Donghong Ji.
\newblock A semantic and syntactic enhanced neural model for financial sentiment analysis.
\newblock \emph{Information Processing \& Management}, 59\penalty0 (4):\penalty0 102943, 2022.

\bibitem[Fei et~al.(2020{\natexlab{a}})Fei, Zhang, Ren, and Ji]{FeiZRJ20}
Hao Fei, Yue Zhang, Yafeng Ren, and Donghong Ji.
\newblock Latent emotion memory for multi-label emotion classification.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial Intelligence}, pages 7692--7699, 2020{\natexlab{a}}.

\bibitem[Tang et~al.(2015)Tang, Qin, Feng, and Liu]{tang2015effective}
Duyu Tang, Bing Qin, Xiaocheng Feng, and Ting Liu.
\newblock Effective lstms for target-dependent sentiment classification.
\newblock \emph{arXiv preprint arXiv:1512.01100}, 2015.

\bibitem[Wu et~al.(2022)Wu, Fei, Li, Zhang, Liu, Teng, and Ji]{Wu0LZLTJ22}
Shengqiong Wu, Hao Fei, Fei Li, Meishan Zhang, Yijiang Liu, Chong Teng, and Donghong Ji.
\newblock Mastering the explicit opinion-role interaction: Syntax-aided neural transition system for unified opinion role labeling.
\newblock In \emph{Proceedings of the Thirty-Sixth {AAAI} Conference on Artificial Intelligence}, pages 11513--11521, 2022.

\bibitem[Ma et~al.(2017)Ma, Li, Zhang, and Wang]{ma2017interactive}
Dehong Ma, Sujian Li, Xiaodong Zhang, and Houfeng Wang.
\newblock Interactive attention networks for aspect-level sentiment classification.
\newblock \emph{arXiv preprint arXiv:1709.00893}, 2017.

\bibitem[Fei et~al.(2020{\natexlab{b}})Fei, Zhang, and Ji]{fei-etal-2020-cross}
Hao Fei, Meishan Zhang, and Donghong Ji.
\newblock Cross-lingual semantic role labeling with high-quality translated training corpus.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics}, pages 7014--7026, 2020{\natexlab{b}}.

\bibitem[Chen et~al.(2017)Chen, Sun, Bing, and Yang]{chen2017recurrent}
Peng Chen, Zhongqian Sun, Lidong Bing, and Wei Yang.
\newblock Recurrent attention network on memory for aspect sentiment analysis.
\newblock In \emph{Proceedings of the 2017 conference on empirical methods in natural language processing}, pages 452--461, 2017.

\bibitem[Fei et~al.(2021{\natexlab{a}})Fei, Zhang, Li, and Ji]{FeiTransiAAAI21}
Hao Fei, Meishan Zhang, Bobo Li, and Donghong Ji.
\newblock End-to-end semantic role labeling with neural transition-based model.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial Intelligence}, pages 12803--12811, 2021{\natexlab{a}}.

\bibitem[Zhang and Qian(2020)]{zhang2020convolution}
Mi~Zhang and Tieyun Qian.
\newblock Convolution over hierarchical syntactic and lexical graphs for aspect level sentiment analysis.
\newblock In \emph{Proceedings of the 2020 conference on empirical methods in natural language processing (EMNLP)}, pages 3540--3549, 2020.

\bibitem[Wu et~al.(2021{\natexlab{b}})Wu, Fei, Ren, Li, Li, and Ji]{WuFRLLJ21}
Shengqiong Wu, Hao Fei, Yafeng Ren, Bobo Li, Fei Li, and Donghong Ji.
\newblock High-order pair-wise aspect and opinion terms extraction with edge-enhanced syntactic graph convolution.
\newblock \emph{{IEEE} {ACM} Trans. Audio Speech Lang. Process.}, 29:\penalty0 2396--2406, 2021{\natexlab{b}}.

\bibitem[Chen et~al.(2020{\natexlab{a}})Chen, Teng, and Zhang]{chen2020inducing}
Chenhua Chen, Zhiyang Teng, and Yue Zhang.
\newblock Inducing target-specific latent structures for aspect sentiment classification.
\newblock In \emph{Proceedings of the 2020 conference on empirical methods in natural language processing (EMNLP)}, pages 5596--5607, 2020{\natexlab{a}}.

\bibitem[Fei et~al.(2022{\natexlab{c}})Fei, Wu, Ren, and Zhang]{FeiMatchStruICML22}
Hao Fei, Shengqiong Wu, Yafeng Ren, and Meishan Zhang.
\newblock Matching structure for dual learning.
\newblock In \emph{Proceedings of the International Conference on Machine Learning, {ICML}}, pages 6373--6391, 2022{\natexlab{c}}.

\bibitem[Huang et~al.(2020)Huang, Sun, Li, Zhang, and Wang]{huang2020syntax}
Lianzhe Huang, Xin Sun, Sujian Li, Linhao Zhang, and Houfeng Wang.
\newblock Syntax-aware graph attention network for aspect-level sentiment classification.
\newblock In \emph{Proceedings of the 28th international conference on computational linguistics}, pages 799--810, 2020.

\bibitem[Fei et~al.(2021{\natexlab{b}})Fei, Li, Li, and Ji]{FeiGraphSynAAAI21}
Hao Fei, Fei Li, Bobo Li, and Donghong Ji.
\newblock Encoder-decoder based unified semantic role labeling with label-aware syntax.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial Intelligence}, pages 12794--12802, 2021{\natexlab{b}}.

\bibitem[Fei et~al.(2023{\natexlab{a}})Fei, Ren, Zhang, and Ji]{9634849}
Hao Fei, Yafeng Ren, Yue Zhang, and Donghong Ji.
\newblock Nonautoregressive encoder-decoder neural framework for end-to-end aspect-based sentiment triplet extraction.
\newblock \emph{IEEE Transactions on Neural Networks and Learning Systems}, 34\penalty0 (9):\penalty0 5544--5556, 2023{\natexlab{a}}.

\bibitem[Hou et~al.(2019)Hou, Huang, Wang, He, and Zhou]{hou2019selective}
Xiaochen Hou, Jing Huang, Guangtao Wang, Xiaodong He, and Bowen Zhou.
\newblock Selective attention based graph convolutional networks for aspect-level sentiment classification.
\newblock \emph{arXiv preprint arXiv:1910.10857}, 2019.

\bibitem[Fei et~al.(2022{\natexlab{d}})Fei, Li, Ren, Zhang, and Ji]{wwwLRZJ22}
Hao Fei, Jingye Li, Yafeng Ren, Meishan Zhang, and Donghong Ji.
\newblock Making decision like human: Joint aspect category sentiment analysis and rating prediction with fine-to-coarse reasoning.
\newblock In \emph{Proceedings of the {ACM} Web Conference 2022, WWW}, pages 3042--3051, 2022{\natexlab{d}}.

\bibitem[Huang and Carley(2019)]{huang2019syntax}
Binxuan Huang and Kathleen~M Carley.
\newblock Syntax-aware aspect level sentiment classification with graph attention networks.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP)}, pages 5472--5480, 2019.

\bibitem[Zhang et~al.(2019)Zhang, Li, and Song]{zhang2019aspect}
Chen Zhang, Qiuchi Li, and Dawei Song.
\newblock Aspect-based sentiment classification with aspect-specific graph convolutional networks.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP)}, pages 4560--4570, 2019.

\bibitem[Sun et~al.(2019)Sun, Zhang, Mensah, Mao, and Liu]{sun2019aspect}
Kai Sun, Richong Zhang, Samuel Mensah, Yongyi Mao, and Xudong Liu.
\newblock Aspect-level sentiment analysis via convolution over dependency tree.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP)}, pages 5683--5692, 2019.

\bibitem[Wang et~al.(2020)Wang, Shen, Yang, Quan, and Wang]{wang2020relational}
Kai Wang, Weizhou Shen, Yunyi Yang, Xiaojun Quan, and Rui Wang.
\newblock Relational graph attention network for aspect-based sentiment analysis.
\newblock \emph{arXiv preprint arXiv:2004.12362}, 2020.

\bibitem[Li et~al.(2022)Li, Fei, Liu, Wu, Zhang, Teng, Ji, and Li]{Li00WZTJL22}
Jingye Li, Hao Fei, Jiang Liu, Shengqiong Wu, Meishan Zhang, Chong Teng, Donghong Ji, and Fei Li.
\newblock Unified named entity recognition as word-word relation classification.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial Intelligence}, pages 10965--10973, 2022.

\bibitem[Fei et~al.(2021{\natexlab{c}})Fei, Ren, Zhang, Ji, and Liang]{fei2020enriching}
Hao Fei, Yafeng Ren, Yue Zhang, Donghong Ji, and Xiaohui Liang.
\newblock Enriching contextualized language model from knowledge graph for biomedical information extraction.
\newblock \emph{Briefings in Bioinformatics}, 22\penalty0 (3), 2021{\natexlab{c}}.

\bibitem[Fei et~al.(2020{\natexlab{c}})Fei, Ren, and Ji]{fei2020boundaries}
Hao Fei, Yafeng Ren, and Donghong Ji.
\newblock Boundaries and edges rethinking: An end-to-end neural model for overlapping entity relation extraction.
\newblock \emph{Information Processing \& Management}, 57\penalty0 (6):\penalty0 102311, 2020{\natexlab{c}}.

\bibitem[Busbridge et~al.(2019)Busbridge, Sherburn, Cavallo, and Hammerla]{busbridge2019relational}
Dan Busbridge, Dane Sherburn, Pietro Cavallo, and Nils~Y Hammerla.
\newblock Relational graph attention networks.
\newblock \emph{arXiv preprint arXiv:1904.05811}, 2019.

\bibitem[Li et~al.(2018{\natexlab{a}})Li, Han, and Wu]{li2018deeper}
Qimai Li, Zhichao Han, and Xiao-Ming Wu.
\newblock Deeper insights into graph convolutional networks for semi-supervised learning.
\newblock In \emph{Thirty-Second AAAI Conference on Artificial Intelligence}, 2018{\natexlab{a}}.

\bibitem[Liu and Zhang(2017)]{liu2017attention}
Jiangming Liu and Yue Zhang.
\newblock Attention modeling for targeted sentiment.
\newblock In \emph{Proceedings of the 15th Conference of the European Chapter of the Association for Computational Linguistics: Volume 2, Short Papers}, pages 572--577, 2017.

\bibitem[Wang et~al.(2018)Wang, Mazumder, Liu, Zhou, and Chang]{wang2018target}
Shuai Wang, Sahisnu Mazumder, Bing Liu, Mianwei Zhou, and Yi~Chang.
\newblock Target-sensitive memory networks for aspect sentiment classification.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, pages 957--967, 2018.

\bibitem[Fan et~al.(2018)Fan, Feng, and Zhao]{fan2018multi}
Feifan Fan, Yansong Feng, and Dongyan Zhao.
\newblock Multi-grained attention network for aspect-level sentiment classification.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing}, pages 3433--3442, 2018.

\bibitem[Zheng and Xia(2018)]{zheng2018left}
Shiliang Zheng and Rui Xia.
\newblock Left-center-right separated neural network for aspect-based sentiment analysis with rotatory attention.
\newblock \emph{arXiv preprint arXiv:1802.00892}, 2018.

\bibitem[Wang and Lu(2018)]{wang2018learning}
Bailin Wang and Wei Lu.
\newblock Learning latent opinions for aspect-level sentiment classification.
\newblock In \emph{Thirty-Second AAAI Conference on Artificial Intelligence}, 2018.

\bibitem[Li et~al.(2018{\natexlab{b}})Li, Liu, and Zhou]{li2018hierarchical}
Lishuang Li, Yang Liu, and AnQiao Zhou.
\newblock Hierarchical attention based position-aware network for aspect-level sentiment analysis.
\newblock In \emph{Proceedings of the 22nd Conference on Computational Natural Language Learning}, pages 181--189, 2018{\natexlab{b}}.

\bibitem[Li et~al.(2018{\natexlab{c}})Li, Bing, Lam, and Shi]{li2018transformation}
Xin Li, Lidong Bing, Wai Lam, and Bei Shi.
\newblock Transformation networks for target-oriented sentiment classification.
\newblock \emph{arXiv preprint arXiv:1805.01086}, 2018{\natexlab{c}}.

\bibitem[Devlin et~al.(2018)Devlin, Chang, Lee, and Toutanova]{devlin2018bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova.
\newblock Bert: Pre-training of deep bidirectional transformers for language understanding.
\newblock \emph{Proceedings of NAACL-HLT 2019}, page pages 4171–4186, 2018.

\bibitem[Song et~al.(2019)Song, Wang, Jiang, Liu, and Rao]{song2019attentional}
Youwei Song, Jiahai Wang, Tao Jiang, Zhiyue Liu, and Yanghui Rao.
\newblock Attentional encoder network for targeted sentiment classification.
\newblock \emph{arXiv preprint arXiv:1902.09314}, 2019.

\bibitem[Xu et~al.(2019)Xu, Liu, Shu, and Yu]{xu2019bert}
Hu~Xu, Bing Liu, Lei Shu, and Philip~S Yu.
\newblock Bert post-training for review reading comprehension and aspect-based sentiment analysis.
\newblock \emph{arXiv preprint arXiv:1904.02232}, 2019.

\bibitem[Rietzler et~al.(2019)Rietzler, Stabinger, Opitz, and Engl]{rietzler2019adapt}
Alexander Rietzler, Sebastian Stabinger, Paul Opitz, and Stefan Engl.
\newblock Adapt or get left behind: Domain adaptation through bert language model finetuning for aspect-target sentiment classification.
\newblock \emph{arXiv preprint arXiv:1908.11860}, 2019.

\bibitem[Fei et~al.(2021{\natexlab{d}})Fei, Ji, Li, Liu, Ren, and Li]{0001JLLRL21}
Hao Fei, Donghong Ji, Bobo Li, Yijiang Liu, Yafeng Ren, and Fei Li.
\newblock Rethinking boundaries: End-to-end recognition of discontinuous mentions with pointer networks.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial Intelligence}, pages 12785--12793, 2021{\natexlab{d}}.

\bibitem[Mukherjee et~al.(2021)Mukherjee, Shetty, Chattopadhyay, Maji, Datta, and Goyal]{mukherjee2021reproducibility}
Rajdeep Mukherjee, Shreyas Shetty, Subrata Chattopadhyay, Subhadeep Maji, Samik Datta, and Pawan Goyal.
\newblock Reproducibility, replicability and beyond: Assessing production readiness of aspect based sentiment analysis in the wild.
\newblock In \emph{Advances in Information Retrieval: 43rd European Conference on IR Research, ECIR 2021, Virtual Event, March 28--April 1, 2021, Proceedings, Part II 43}, pages 92--106. Springer, 2021.

\bibitem[Fei et~al.(2023{\natexlab{b}})Fei, Chua, Li, Ji, Zhang, and Ren]{FeiCLJZR23}
Hao Fei, Tat{-}Seng Chua, Chenliang Li, Donghong Ji, Meishan Zhang, and Yafeng Ren.
\newblock On the robustness of aspect-based sentiment analysis: Rethinking model, data, and training.
\newblock \emph{{ACM} Transactions on Information Systems}, 41\penalty0 (2):\penalty0 50:1--50:32, 2023{\natexlab{b}}.

\bibitem[Zhuang et~al.(2023)Zhuang, Fei, and Hu]{ZhuangFH23}
Ling Zhuang, Hao Fei, and Po~Hu.
\newblock Knowledge-enhanced event relation extraction via event ontology prompt.
\newblock \emph{Inf. Fusion}, 100:\penalty0 101919, 2023.

\bibitem[Fei et~al.(2021{\natexlab{e}})Fei, Wu, Ren, Li, and Ji]{FeiWRLJ21}
Hao Fei, Shengqiong Wu, Yafeng Ren, Fei Li, and Donghong Ji.
\newblock Better combine them together! integrating syntactic constituency and dependency representations for semantic role labeling.
\newblock In \emph{Findings of the Association for Computational Linguistics: {ACL/IJCNLP} 2021}, pages 549--559, 2021{\natexlab{e}}.

\bibitem[Chen et~al.(2020{\natexlab{b}})Chen, Sun, Wang, Li, Si, Zhang, and Zhou]{chen2020aspect}
Xiao Chen, Changlong Sun, Jingjing Wang, Shoushan Li, Luo Si, Min Zhang, and Guodong Zhou.
\newblock Aspect sentiment classification with document-level sentiment preference modeling.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics}, pages 3667--3677, 2020{\natexlab{b}}.

\bibitem[Li et~al.(2021)Li, Xu, Li, Fei, Ren, and Ji]{li-etal-2021-mrn}
Jingye Li, Kang Xu, Fei Li, Hao Fei, Yafeng Ren, and Donghong Ji.
\newblock {MRN}: A locally and globally mention-based reasoning network for document-level relation extraction.
\newblock In \emph{Findings of the Association for Computational Linguistics: ACL-IJCNLP 2021}, pages 1359--1370, 2021.

\bibitem[Liu et~al.(2023)Liu, Fei, Li, Li, Li, Zhao, Teng, and Ji]{TKDP2306}
Jiang Liu, Hao Fei, Fei Li, Jingye Li, Bobo Li, Liang Zhao, Chong Teng, and Donghong Ji.
\newblock {TKDP:} threefold knowledge-enriched deep prompt tuning for few-shot named entity recognition.
\newblock \emph{CoRR}, abs/2306.03974, 2023.

\bibitem[Pontiki et~al.(2016)Pontiki, Galanis, Papageorgiou, Androutsopoulos, Manandhar, AL-Smadi, Al-Ayyoub, Zhao, Qin, De~Clercq, et~al.]{pontiki2016semeval}
Maria Pontiki, Dimitris Galanis, Haris Papageorgiou, Ion Androutsopoulos, Suresh Manandhar, Mohammed AL-Smadi, Mahmoud Al-Ayyoub, Yanyan Zhao, Bing Qin, Orph{\'e}e De~Clercq, et~al.
\newblock Semeval-2016 task 5: Aspect based sentiment analysis.
\newblock In \emph{ProWorkshop on Semantic Evaluation (SemEval-2016)}, pages 19--30. Association for Computational Linguistics, 2016.

\bibitem[Fei et~al.(2022{\natexlab{e}})Fei, Li, Li, Wu, Li, and Ji]{Feiijcai22UABSA}
Hao Fei, Fei Li, Chenliang Li, Shengqiong Wu, Jingye Li, and Donghong Ji.
\newblock Inheriting the wisdom of predecessors: A multiplex cascade framework for unified aspect-based sentiment analysis.
\newblock In \emph{Proceedings of the Thirty-First International Joint Conference on Artificial Intelligence, {IJCAI}}, pages 4096--4103, 2022{\natexlab{e}}.

\bibitem[Wang et~al.(2022)Wang, Li, Fei, Li, Wu, Su, Shi, Ji, and Cai]{wang-etal-2022-entity}
Fengqi Wang, Fei Li, Hao Fei, Jingye Li, Shengqiong Wu, Fangfang Su, Wenxuan Shi, Donghong Ji, and Bo~Cai.
\newblock Entity-centered cross-document relation extraction.
\newblock In \emph{Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing}, pages 9871--9881, 2022.

\bibitem[Fei et~al.(2020{\natexlab{d}})Fei, Ren, and Ji]{FeiRJ20}
Hao Fei, Yafeng Ren, and Donghong Ji.
\newblock Retrofitting structure-aware transformer language model for end tasks.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing}, pages 2151--2161, 2020{\natexlab{d}}.

\bibitem[Jia et~al.(2021)Jia, Wang, Zan, and Xie]{jia2021syntactic}
Yuxiang Jia, Yadong Wang, Hongying Zan, and Qi~Xie.
\newblock Syntactic information and multiple semantic segments for aspect-based sentiment classification.
\newblock \emph{International Journal of Asian Language Processing}, 31\penalty0 (03n04):\penalty0 2250006, 2021.

\bibitem[Cao et~al.(2022)Cao, Li, Su, Li, Fei, Wu, Li, Zhao, and Ji]{cao-etal-2022-oneee}
Hu~Cao, Jingye Li, Fangfang Su, Fei Li, Hao Fei, Shengqiong Wu, Bobo Li, Liang Zhao, and Donghong Ji.
\newblock {O}ne{EE}: A one-stage framework for fast overlapping and nested event extraction.
\newblock In \emph{Proceedings of the 29th International Conference on Computational Linguistics}, pages 1953--1964, 2022.

\bibitem[He et~al.(2018)He, Lee, Ng, and Dahlmeier]{he2018effective}
Ruidan He, Wee~Sun Lee, Hwee~Tou Ng, and Daniel Dahlmeier.
\newblock Effective attention modeling for aspect-level sentiment classification.
\newblock In \emph{Proceedings of the 27th International Conference on Computational Linguistics}, pages 1121--1131, 2018.

\bibitem[Nguyen and Shirai(2015)]{nguyen2015phrasernn}
Thien~Hai Nguyen and Kiyoaki Shirai.
\newblock Phrasernn: Phrase recursive neural network for aspect-based sentiment analysis.
\newblock In \emph{Proceedings of the 2015 Conference on Empirical Methods in Natural Language Processing}, pages 2509--2514, 2015.

\bibitem[Fei et~al.(2020{\natexlab{e}})Fei, Ren, and Ji]{fei-etal-2020-improving}
Hao Fei, Yafeng Ren, and Donghong Ji.
\newblock Improving text understanding via deep syntax-semantics communication.
\newblock In \emph{Findings of the Association for Computational Linguistics: EMNLP 2020}, pages 84--93, 2020{\natexlab{e}}.

\bibitem[Dozat and Manning(2016)]{dozat2016deep}
Timothy Dozat and Christopher~D Manning.
\newblock Deep biaffine attention for neural dependency parsing.
\newblock \emph{arXiv preprint arXiv:1611.01734}, 2016.

\bibitem[Wu et~al.(2023)Wu, Fei, Qu, Ji, and Chua]{wu2023nextgpt}
Shengqiong Wu, Hao Fei, Leigang Qu, Wei Ji, and Tat-Seng Chua.
\newblock Next-gpt: Any-to-any multimodal llm, 2023.

\bibitem[Veli{\v{c}}kovi{\'c} et~al.(2017)Veli{\v{c}}kovi{\'c}, Cucurull, Casanova, Romero, Lio, and Bengio]{velivckovic2017graph}
Petar Veli{\v{c}}kovi{\'c}, Guillem Cucurull, Arantxa Casanova, Adriana Romero, Pietro Lio, and Yoshua Bengio.
\newblock Graph attention networks.
\newblock \emph{arXiv preprint arXiv:1710.10903}, 2017.

\bibitem[Tang et~al.(2020)Tang, Ji, Li, and Zhou]{tang-etal-2020-dependency}
Hao Tang, Donghong Ji, Chenliang Li, and Qiji Zhou.
\newblock Dependency graph enhanced dual-transformer structure for aspect-based sentiment classification.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics}, pages 6578--6588, Online, July 2020. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2020.acl-main.588}.
\newblock URL \url{https://www.aclweb.org/anthology/2020.acl-main.588}.

\bibitem[Manning et~al.(2014)Manning, Surdeanu, Bauer, Finkel, Bethard, and McClosky]{manning2014stanford}
Christopher Manning, Mihai Surdeanu, John Bauer, Jenny Finkel, Steven Bethard, and David McClosky.
\newblock The stanford corenlp natural language processing toolkit.
\newblock In \emph{Proceedings of 52nd annual meeting of the association for computational linguistics: system demonstrations}, pages 55--60, 2014.

\bibitem[Qi et~al.(2020)Qi, Zhang, Zhang, Bolton, and Manning]{qi2020stanza}
Peng Qi, Yuhao Zhang, Yuhui Zhang, Jason Bolton, and Christopher~D. Manning.
\newblock Stanza: A {Python} natural language processing toolkit for many human languages.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics: System Demonstrations}, 2020.
\newblock URL \url{https://nlp.stanford.edu/pubs/qi2020stanza.pdf}.

\bibitem[Kitaev and Klein(2018)]{Kitaev-2018-SelfAttentive}
Nikita Kitaev and Dan Klein.
\newblock Constituency parsing with a self-attentive encoder.
\newblock In \emph{Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers)}, Melbourne, Australia, July 2018. Association for Computational Linguistics.

\bibitem[Paszke et~al.(2019)Paszke, Gross, Massa, Lerer, Bradbury, Chanan, Killeen, Lin, Gimelshein, Antiga, Desmaison, Kopf, Yang, DeVito, Raison, Tejani, Chilamkurthy, Steiner, Fang, Bai, and Chintala]{NEURIPS2019_9015}
Adam Paszke, Sam Gross, Francisco Massa, Adam Lerer, James Bradbury, Gregory Chanan, Trevor Killeen, Zeming Lin, Natalia Gimelshein, Luca Antiga, Alban Desmaison, Andreas Kopf, Edward Yang, Zachary DeVito, Martin Raison, Alykhan Tejani, Sasank Chilamkurthy, Benoit Steiner, Lu~Fang, Junjie Bai, and Soumith Chintala.
\newblock Pytorch: An imperative style, high-performance deep learning library.
\newblock In H.~Wallach, H.~Larochelle, A.~Beygelzimer, F.~d\textquotesingle Alch\'{e}-Buc, E.~Fox, and R.~Garnett, editors, \emph{Advances in Neural Information Processing Systems 32}, pages 8024--8035. Curran Associates, Inc., 2019.
\newblock URL \url{http://papers.neurips.cc/paper/9015-pytorch-an-imperative-style-high-performance-deep-learning-library.pdf}.

\bibitem[Wang et~al.(2019)Wang, Yu, Zheng, Gan, Gai, Ye, Li, Zhou, Huang, Ma, Huang, Guo, Zhang, Lin, Zhao, Li, Smola, and Zhang]{wang2019dgl}
Minjie Wang, Lingfan Yu, Da~Zheng, Quan~Gan Gan, Yu~Gai, Zihao Ye, Mufei Li, Jinjing Zhou, Qi~Huang, Chao Ma, Ziyue Huang, Qipeng Guo, Hao Zhang, Haibin Lin, Junbo Zhao, Jinyang Li, Alexander Smola, and Zheng Zhang.
\newblock Deep graph library: Towards efficient and scalable deep learning on graphs.
\newblock 2019.

\bibitem[Srivastava et~al.(2014)Srivastava, Hinton, Krizhevsky, Sutskever, and Salakhutdinov]{srivastava2014dropout}
Nitish Srivastava, Geoffrey Hinton, Alex Krizhevsky, Ilya Sutskever, and Ruslan Salakhutdinov.
\newblock Dropout: a simple way to prevent neural networks from overfitting.
\newblock \emph{The Journal of Machine Learning Research}, 15\penalty0 (1):\penalty0 1929--1958, 2014.

\end{thebibliography}

\documentclass{article}

\usepackage{myarxiv}

\usepackage[utf8]{inputenc} % allow utf-8 input
\usepackage[T1]{fontenc}    % use 8-bit T1 fonts
\usepackage{hyperref}       % hyperlinks
\usepackage{url}            % simple URL typesetting
\usepackage{booktabs}       % professional-quality tables
\usepackage{amsfonts}       % blackboard math symbols
\usepackage{nicefrac}       % compact symbols for 1/2, etc.
\usepackage{microtype}      % microtypography
\usepackage{lipsum}		% Can be removed after putting your text content
\usepackage{graphicx}
\usepackage[numbers]{natbib}
\usepackage{doi}
\usepackage{latexsym}
\usepackage{stmaryrd}
\usepackage{amsmath,amssymb,amsfonts}
\usepackage{multicol,lipsum}
\usepackage{subcaption}
\usepackage{amsmath,amssymb}

\usepackage{threeparttable}

\usepackage{enumitem}

\usepackage{multirow}

\title{Syntactic Fusion: Enhancing Aspect-Level Sentiment Analysis Through Multi-Tree Graph Integration}

\author{ 
Jane Sunny\\
University of Charleston\\
\And
Tom Padraig\\
University of Charleston\\
\And
Roggie Terry\\
University of Charleston\\
\And
Woods Ali\\
University of Charleston\\
}

\renewcommand{\shorttitle}{Syntactic Fusion: Enhancing Aspect-Level Sentiment Analysis Through Multi-Tree Graph Integration}

\newcommand{\methodname}{SynthFusion}

\hypersetup{
pdftitle={aaa},
pdfsubject={aaa},
pdfauthor={aaa},
pdfkeywords={aaa},
}

\begin{document}
\maketitle

\begin{abstract}
Recent progress in aspect-level sentiment classification has been propelled by the incorporation of graph neural networks (GNNs) leveraging syntactic structures, particularly dependency trees. Nevertheless, the performance of these models is often hampered by the innate inaccuracies of parsing algorithms. To mitigate this challenge, we introduce SynthFusion, an innovative graph ensemble method that amalgamates predictions from multiple parsers. This strategy blends diverse dependency relations prior to the application of GNNs, enhancing robustness against parsing errors while avoiding extra computational burdens. SynthFusion circumvents the pitfalls of overparameterization and diminishes the risk of overfitting, prevalent in models with stacked GNN layers, by optimizing graph connectivity. Our empirical evaluations on the SemEval14 and Twitter14 datasets affirm that SynthFusion not only outshines models reliant on single dependency trees but also eclipses alternative ensemble techniques, achieving this without an escalation in model complexity.
\end{abstract}

\keywords{Sentiment Analysis \and Syntax Feature Modeling}

\section{Introduction}

Aspect-level sentiment classification is a nuanced task within sentiment analysis, focused on discerning the sentiment (positive, negative, or neutral) towards specific aspects in textual content. For instance, in the sentence ``\textit{The exterior, unlike the food, is unwelcoming.}'', the sentiments towards ``exterior'' and ``food'' are negative and positive respectively \citep{nazir2020issues,FeiLasuieNIPS22,kiritchenkodetecting,dong2014adaptive}.
This task is particularly useful in parsing online reviews or aiding purchasing decisions on e-commerce platforms.

Recent studies highlight the effectiveness of syntactic structures, such as dependency trees, in capturing complex syntactic relationships often overlooked in superficial text analysis~\cite{zhang2018graph}. \nocite{FeiDiaREIJCAI22,Wu0RJL21,xiang2022semantic,FeiZRJ20,tang2015effective,Wu0LZLTJ22,ma2017interactive,fei-etal-2020-cross,chen2017recurrent,FeiTransiAAAI21,zhang2020convolution,WuFRLLJ21,chen2020inducing,FeiMatchStruICML22,huang2020syntax,FeiGraphSynAAAI21,9634849,hou2019selective,wwwLRZJ22}
Techniques employing graph neural networks (GNN) over dependency trees for aspect-level sentiment classification have shown promising results, linking aspect terms with relevant opinions for robust sentiment analysis~\cite{huang2019syntax,zhang2019aspect,sun2019aspect,wang2020relational}. However, these methods are not immune to the errors in parsing algorithms~\cite{wang2020relational}. For example, a misparsed sentence could lead to incorrect sentiment association for the term ``food''.

Acknowledging the limitations of current dependency parsers, especially in scenarios diverging from their training domain, we propose \methodname{}, a graph ensemble technique to mitigate parsing inaccuracies. This approach is grounded in the observation that different parsers, particularly those with distinct biases, often make unique errors \citep{Li00WZTJL22,fei2020enriching,dong2014adaptive,tang2015effective,fei2020boundaries}. By combining dependency trees from various parsers before implementing GNNs, our technique allows the model to concurrently assess multiple parsing hypotheses, thereby enhancing its decision-making process.

\methodname{} fuses edges from all provided dependency trees to create a comprehensive ensemble graph for GNN application. This exposes the model to diverse graph structures, empowering it to discern and prioritize more relevant edges for the given task. To maintain the syntactic relationships from the original trees, we categorize edges into parent-to-children and children-to-parent types, applying relational graph attention networks (RGAT)~\cite{busbridge2019relational} to the ensemble graph.

Our approach presents several advantages. Firstly, \methodname{}'s integration of multiple dependency trees allows GNNs to navigate through various parsing scenarios, enhancing the model's resilience to parsing errors. Moreover, this is achieved without incurring additional computational costs, as the GNNs operate on a singular graph. Additionally, by introducing more edges and reducing graph diameter, \methodname{} prevents GNNs from overfitting and limits the need for over-parameterization. This facilitates the use of shallower GNN models, thereby circumventing the common issue of over-smoothing in deeper networks~\cite{li2018deeper}.

In summary, our work contributes the following:
\begin{itemize}
    \item Introducing \methodname{}, an innovative technique that amalgamates dependency trees from diverse parsers to bolster model resilience against parsing errors. This ensemble graph approach allows the model to learn from noisy data and discern the most relevant edges without extra computational demands.
    
    \item Maintaining syntactic dependency integrity from the original trees by uniquely parameterizing edges, thereby enhancing the RGAT model's performance on the ensemble graph.
    
    \item Demonstrating that our \methodname{} RGAT model outshines recent leading models on three benchmark datasets (Laptop and Restaurant reviews from SemEval 2014 and the ACL 14 Twitter dataset), surpassing both single-parse models and other ensemble strategies.
\end{itemize}

\section{Related Work}

The field of aspect-level sentiment classification has witnessed significant advancements, with many studies emphasizing the application of attention mechanisms to sequential models. These attention mechanisms include co-attention, self-attention, and hierarchical attention, which have been effectively incorporated into recurrent neural networks (RNNs) for enhanced sentiment analysis~\cite{tang2015effective,liu2017attention,wang2018target,fan2018multi,chen2017recurrent,zheng2018left,wang2018learning,li2018hierarchical,li2018transformation}. These approaches leverage the sequential nature of language to better capture the nuances of sentiment associated with specific aspects in text.

Additionally, the advent of pretrained transformer language models, notably BERT~\cite{devlin2018bert}, has introduced a new dimension to this task. Unlike traditional models that rely on engineered features or complex architectures, BERT and similar models utilize deep bidirectional representations learned from vast amounts of unlabeled text. This has enabled more direct and effective handling of word sequences in sentiment classification, as demonstrated in various studies~\cite{song2019attentional,xu2019bert,rietzler2019adapt}.

Simultaneously, there's growing interest in integrating syntactic information into aspect-level sentiment analysis. \nocite{0001JLLRL21,mukherjee2021reproducibility,FeiCLJZR23,ZhuangFH23,zhang2019aspect,FeiWRLJ21,chen2020aspect,li-etal-2021-mrn,TKDP2306,pontiki2016semeval,Feiijcai22UABSA,wang-etal-2022-entity,FeiRJ20,jia2021syntactic,cao-etal-2022-oneee} Researchers have explored the use of both dependency and constituency trees to capture syntactic relationships within sentences. Dependency trees, which represent grammatical structures, have been particularly instrumental, with several studies employing them to improve sentiment classification models~\cite{dong2014adaptive, he2018effective}. Constituency trees, which represent sentence structure based on phrase constituency, have also been used to enhance sentiment analysis, as seen in works like~\cite{nguyen2015phrasernn}.

Recent research has further explored the potential of Graph Neural Networks (GNNs) in processing syntactic structures, leading to the development of robust dependency-based models. These models, which operate either directly on dependency trees or on specially reshaped trees centering around aspect terms, have shown promise in enhancing sentiment classification accuracy~\cite{huang2019syntax, zhang2019aspect, sun2019aspect, wang2020relational}. 
In a novel approach, \citet{fei-etal-2020-improving} combined GNNs and BERT models, reporting improved performance through a joint learning mechanism facilitated by mutual biaffine attention \citep{dozat2016deep}.

However, a persistent challenge for these dependency-based models is their vulnerability to parsing errors. This is primarily because they often rely on a single parser's output, which may not be error-free. Additionally, most of these models treat dependency trees as undirected graphs, thereby overlooking the rich syntactic relationships that these trees encode. As such, there remains an untapped potential in considering the directional nature of these trees to better understand and utilize the syntactic relationships between words in a sentence for sentiment analysis \citep{nazir2020issues,FeiLasuieNIPS22,kiritchenkodetecting,wu2023nextgpt}.

\section{Framework}

Our focus lies in discerning the sentiment polarity of an aspect term within a sentence. For a sentence with $n$ words $\{w_1,w_2,\ldots,w_{\tau}, \ldots,w_{\tau+t}, \ldots,w_n\}$, containing the aspect term $\{w_{\tau}, w_{\tau+1},\ldots,w_{\tau+t-1}\}$, the objective is to categorize the sentiment directed at this term as either positive, negative, or neutral. While applying Graph Neural Networks (GNNs) on dependency trees has proven effective for this task, it remains susceptible to parsing inaccuracies. To counter this, we introduce \methodname{}, a technique that integrates multiple dependency trees, enhancing error resilience. This section details the \methodname{} graph ensemble method and its application in a GNN model for aspect-level sentiment analysis.

\subsection{\methodname{} over Multiple Dependency Trees}

\methodname{} is designed to fortify graph neural networks against parse errors by amalgamating various parses into a unified ensemble graph. For a given sentence $\{w_1,w_2,\ldots,w_n\}$ and $M$ distinct dependency parses $G_1, \ldots, G_M$, \methodname{} creates a composite graph $G$:
\begin{flalign}
 G &= (V, \{e | e = (w_i, w_j) \in \bigcup_{m=1}^M E_m\})
\end{flalign}
Here, $V$ represents the shared nodes set across all graphs, assuming parsers have identical tokenization, and $E_m$ denotes the edges in $G_m$. The ensemble graph $G$, includes all directed edges from each dependency tree.

This approach allows GNN models to operate on $G$ just as they would on a single dependency tree, without incurring extra computational costs common in traditional ensemble methods. Importantly, $G$ encompasses more edges from the accurate dependency tree, thus reinforcing the GNN's robustness against individual parser errors. Additionally, the increased connectivity in $G$ reduces the graph's diameter, enabling shallower GNN models to learn effectively, thereby minimizing the risk of over-parameterization and overfitting.

\subsection{RGAT over Ensemble Graph}

To derive node representations from the ensemble graph, we employ Graph Attention Networks (GAT)~\cite[GAT;][]{velivckovic2017graph}. Each node's hidden representation in GAT is formulated by attending over its neighboring nodes using a multi-head self-attention mechanism. The representation of word $i$ at the $l$-th GAT layer is:
\begin{equation}
{h}^{(l)}_i =\mathbin\Vert_{k=1}^{K}\sigma(\sum_{j \in N_i}\alpha_{ij}^k{W}^k{h}^{(l-1,k)}_i)
\end{equation}
where $K$ is the number of attention heads, $N_i$ denotes node $i$’s neighborhood, and $\mathbin\Vert$ signifies concatenation. $\bf{W}^k \in \mathbb{R}^{d_B \times d_h}$ are the learnable weights in GAT and $\sigma$ represents the $\mathrm{ReLU}$ activation function. The attention score $\alpha_{ij}^k$ quantifies the importance of node $j$ to node $i$ under head $k$.

\textbf{Edge Types.}
In applying GAT to ensemble graphs, we introduce reciprocal edges for each dependency tree edge, categorizing them as parent-to-children and children-to-parent. This retains the original syntactic relationships between words. Additionally, each graph node receives a self-loop, distinct from dependency edges, creating a third edge type.

We implement Relational GAT (RGAT) to consider these edge types. The attention score computation incorporates edge-type information:
\begin{align}
\alpha_{ij}&=\frac{\exp(\sigma(\bf{a}\bf{W}(\bf{h}_i\|\bf{h}_j)+\bf{a}_e\bf{e}_{ij}))}{\sum_{v \in N_i}\exp(\sigma(\bf{a}\bf{W}(\bf{h}_i\|\bf{h}_v)+\bf{a}_e\bf{e}_{iv}))} \label{eq2}
\end{align}
where $\bf{e}_{ij}$ is the edge type's representation connecting nodes $i$ and $j$. $\bf{a} \in \mathbb{R}^{d_h}$, $\bf{W} \in \mathbb{R}^{d_h \times 2d_h}$, and $\bf{a}_e \in \mathbb{R}^{d_h}$ are learnable matrices.

\subsection{Sentiment Classification}
For sentiment classification, we extract hidden representations of aspect term nodes from the final RGAT layer and apply average pooling to get $\bf{h}_t \in \mathbb{R}^{d_{h}}$. These are then processed through a two-layer MLP to generate the classification scores $\hat{\bf{y}}_s$:
\begin{flalign}
\hat{\bf{y}}_s&=\mathrm{softmax}({\bf{W}_{2}\mathrm{ReLU}({\bf{W}_{1}\bf{h}_{t}})})
\end{flalign}
where $\bf{W}_{2} \in \mathbb{R}^{C \times d_{out}}$ and $\bf{W}_{1} \in \mathbb{R}^{d_{out} \times d_{h}}$ are the weight matrices, with $C$ representing the sentiment class count. We aim to minimize the standard cross-entropy loss function and apply weight decay to the model parameters.

\subsection{RGAT Input}

Initial word node features for RGAT are derived from a BERT encoder, complemented by positional embeddings.

\textbf{BERT Encoder.}
We utilize the BERT base model for initial word representations. The input format is ``[CLS] + sentence + [SEP] + term + [SEP]'', which BERT processes to produce term-centric representations. Subword-based outputs from BERT are averaged to match word-based RGAT inputs, forming the raw RGAT input $\bf{X}$.

\textbf{Positional Encoding.}
Given the task's sensitivity to word position, particularly in sentences with multiple aspects, we incorporate explicit positional encoding. This complements BERT's inherent positional awareness, which may be diluted through Transformer layers. A trainable position embedding matrix is added to $\bf{X}$ before feeding it into RGAT.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\begin{table}[!t]
\centering
\small
% \begin{adjustbox}{max width=0.5\textwidth}
\begin{tabular}{ c c c c c c c }
\toprule
\multirow{2} {*}{Dataset} & \multicolumn{2}{c}{Positive} & \multicolumn{2}{c}{Neutral} & \multicolumn{2}{c}{Negative}\\
\cline{2-7}
& Train & Test & Train & Test & Train & Test\\
\hline
Laptop &  \phantom{0}987 & 341 & \phantom{0}460 & 169 & \phantom{0}866 & 128\\
Restaurant & 2164 & 728 & \phantom{0}633 & 196 & \phantom{0}805 & 196\\
Twitter & 1561 & 173 & 3127 & 346 & 1560 & 173 \\
\bottomrule
\end{tabular}
% \end{adjustbox}
\caption{Statistics of {the three benchmark datasets used in our experiments}.}
\label{table:data}
\end{table}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\begin{table}[!t]
\centering
\resizebox{1.\columnwidth}{!}{
\begin{threeparttable}
\begin{tabular}{c c c c c c c c}
\toprule
\multirow{2} {*}{Category}  &\multirow{2} {*}{Model}  & \multicolumn{2}{c}{14Rest} & \multicolumn{2}{c}{14Lap} & \multicolumn{2}{c}{Twitter} \\
\cline{3-8}
 & & Acc & Macro-F1 & Acc & Macro-F1 & Acc & Macro-F1 \\
\midrule
\multirow{2} {*}{BERT}&BERT-SPC~\cite{song2019attentional} 
& 84.46 & 76.98 & 78.99 & 75.03 & 73.55 & 72.14 \\
&AEN-BERT~\cite{song2019attentional} 
& 83.12 & 73.76 & 79.93 & 76.31 & 74.71 & 73.13  \\
\hline
BERT+DT\tnote{$\star$} &DGEDT-BERT~\cite{tang-etal-2020-dependency}  & 86.3 & 80.0 & 79.8 & 75.6 & 77.9 & 75.4\\
\hline
BERT+RDT{$^\diamond$} & R-GAT+BERT~\cite{wang2020relational} & 86.60  & 81.35  & 78.21 & 74.07 & 76.15 & 74.88 \\
\hline
% \cline{2-8}
Ours &{\textbf{\methodname{}}} & \textbf{87.32} & \textbf{81.95} &  \textbf{81.35} &  \textbf{78.65} & \textbf{78.18} & \textbf{76.52}\\
\bottomrule
\end{tabular}
% \begin{tablenotes}
%      \item 
% \end{tablenotes}
\end{threeparttable}
}
\caption{Comparison of our \methodname{} model to different published numbers on three datasets, with the same setup of train and test data -- no dev data. $\star$ DT: Dependency Tree; $\diamond$ RDT: Reshaped Dependency Tree. The bold text indicates the best results.}
\label{table:bestresults}
\end{table}

\section{Experiment}

\subsection{Settings} 
Our experiments are conducted on three pivotal datasets: Restaurant and Laptop reviews from the SemEval 2014 Task 4 {(14Rest and 14Lap)}\footnote{https://alt.qcri.org/semeval2014/task4/}, and the extensively used ACL 14 Twitter dataset {(Twitter)}~\cite{dong2014adaptive}. Instances with "conflict" sentiments are excluded to streamline the datasets. Detailed statistics of these datasets are presented in Table \ref{table:data}. We assess our model's performance using accuracy and macro F1 scores, in line with standard practices.

For dependency parsing, we utilize Stanford CoreNLP~\cite{manning2014stanford}, Stanza~\cite{qi2020stanza}, and the Berkeley neural parser~\cite{Kitaev-2018-SelfAttentive}. The Berkeley parser generates constituency parses, which we convert to dependency parses using CoreNLP for uniformity.

% \textbf{Comparative Analysis with Baselines.} 
Our \methodname{} model is benchmarked against several established models on these datasets. We include BERT-SPC~\cite{song2019attentional}, which inputs sentence-term pairs into BERT; AEN-BERT~\cite{song2019attentional} utilizes BERT as an encoder with additional attention layers; DGEDT-BERT~\cite{tang-etal-2020-dependency} integrates Transformer representations with GNN models over dependency trees; R-GAT+BERT~\cite{wang2020relational} reshapes dependency trees around aspect terms and encodes them using RGAT. Our model's performance is reported with identical data splits as these models, excluding any development set.

We also implement several baseline models for an in-depth analysis, including:
\begin{enumerate}[itemsep=1pt,leftmargin=10pt]
    \item \textit{BERT-baseline}, which processes sentence-term pairs using the BERT-base encoder followed by a classifier.
    \item \textit{GAT-baseline with Stanza}, using a vanilla GAT model over Stanza's dependency tree, with BERT encoder outputs as initial node features.
    \item \textit{RGAT over single dependency trees}, applying RGAT models with distinct edge types to different dependency trees parsed by CoreNLP, Stanza, and Berkeley parsers, and leveraging BERT encoder outputs enhanced with positional embeddings.
    \item Ensemble models, including \textit{Label-Ensemble}, aggregating outputs from models trained on different parses, and \textit{Feature-Ensemble}, combining features from RGAT models each dedicated to a different parse.
\end{enumerate}

% \textbf{Model Training and Optimization.} 
The models are implemented using Pytorch~\cite{NEURIPS2019_9015} and GAT based on Deep Graph Library~\cite{wang2019dgl}. We set the learning rate at $10^{-5}$ and batch size at 4. Model hyperparameters, including hidden dimensions, attention heads, and GAT/RGAT layers, are optimized using a development set comprising 5\% of the training data. The best model is identified based on the performance on the development set, employing dropout~\cite{srivastava2014dropout} and L2 regularization for generalization. The models are trained for a maximum of 5 epochs.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\begin{table}
\centering
\resizebox{1.0\columnwidth}{!}{
\begin{tabular}{c c c c c c c}
\toprule
\multirow{2} {*}{Model}  & \multicolumn{2}{c}{14Rest} & \multicolumn{2}{c}{14Lap} & \multicolumn{2}{c}{Twitter} \\
\cline{2-7}
  & Acc & Macro-F1 & Acc & Macro-F1 & Acc & Macro-F1 \\
\midrule
 BERT-baseline & 83.43 $\pm$ 0.52 & 74.94 $\pm$ 1.37 & 77.34 $\pm$ 0.90 & 72.77 $\pm$ 1.96 & 73.47 $\pm$ 0.89 & 72.63 $\pm$ 0.82\\
{GAT-baseline with Stanza} & 84.29 $\pm$ 0.30 & 75.75 $\pm$ 1.11 & 77.84 $\pm$ 0.27 & 74.0 $\pm$ 0.55 & 73.82 $\pm$ 0.70 & 72.72 $\pm$ 0.42 \\
\midrule
{RGAT with Stanza} & 84.53 $\pm$ 0.66 & 77.29 $\pm$ 1.42 & 77.99 $\pm$ 0.62 & 74.35 $\pm$ 0.60 & 73.99 $\pm$ 0.48 & 72.76 $\pm$ 0.33 \\
{RGAT with Berkeley} & 84.41 $\pm$ 0.86 & 76.63 $\pm$ 1.38 & 78.09 $\pm$ 1.24 & 73.65 $\pm$ 1.76 & 74.07 $\pm$ 0.67 & 72.65 $\pm$ 0.74 \\
{RGAT with CoreNLP} & 83.86 $\pm$ 0.32 & 76.03 $\pm$ 0.88 & 78.12 $\pm$ 1.02 & 73.86 $\pm$ 1.72 & 73.96 $\pm$ 0.93 & 72.83 $\pm$ 0.95 \\
\midrule
{Label-Ensemble} & 84.68 $\pm$ 0.95 & 77.21 $\pm$ 1.54 & 78.40 $\pm$ 1.51 & 74.36 $\pm$2.45 & 74.59 $\pm$ 0.46 & 73.51 $\pm$ 0.43 \\
{Feature-Ensemble} & 84.64 $\pm$ 0.77 & 77.06 $\pm$ 1.45 & 78.68 $\pm$ 0.69 & 74.80 $\pm$ 0.92 & 74.62 $\pm$ 0.76 & 73.61 $\pm$ 0.73 \\
\midrule
{\textbf{\methodname{}}} & \textbf{85.16} $\pm$ 0.53 &  \textbf{77.91} $\pm$ 0.87 &  \textbf{80.00} $\pm$ 0.63 &  \textbf{76.50} $\pm$ 0.64 &  \textbf{74.74} $\pm$ 0.93 &  \textbf{73.66} $\pm$ 0.88 \\
\bottomrule
\end{tabular}
}
\caption{Comparison of our \methodname{} model to different baselines on three datasets, with 5\% dev data set aside. The bold text indicates the best results.}
\label{table:results}
\end{table}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\begin{table}[!h]
\centering
\resizebox{1\columnwidth}{!}{
\begin{threeparttable}
\begin{tabular}{l c c c c c c}
\toprule
\multirow{2} {*}{Model}  & \multicolumn{2}{c}{14Rest} & \multicolumn{2}{c}{14Lap} & \multicolumn{2}{c}{Twitter}\\
\cline{2-7}
& Acc & Macro-F1 & Acc & Macro-F1 & Acc & Macro-F1\\
\midrule
\textbf{\methodname{}} & \textbf{85.16} $\pm$ 0.53 &  \textbf{77.91} $\pm$ 0.87 &  \textbf{80.00} $\pm$ 0.63 &  \textbf{76.50} $\pm$ 0.64 &  \textbf{74.74} $\pm$ 0.93 &  \textbf{73.66} $\pm$ 0.88 \\
\midrule
- Edge type & 84.25 $\pm$ 0.59 & 76.15 $\pm$ 1.24 & 78.65 $\pm$ 0.51 & 74.76 $\pm$ 0.71 & 74.37 $\pm$ 1.08 & 73.25 $\pm$ 0.85 \\
- Position & 84.36 $\pm$ 0.36 & 75.92 $\pm$ 1.18 & 78.37 $\pm$ 0.31 & 74.51 $\pm$ 0.48 & 74.28 $\pm$ 1.39 & 73.34 $\pm$ 1.35\\
- (Edge type + Position) & 84.16 $\pm$ 0.31 & 75.38 $\pm$ 0.69 & 78.09 $\pm$ 0.27 & 74.29 $\pm$ 0.64 & 73.41 $\pm$ 0.63 & 72.52 $\pm$ 0.62\\
\midrule
{Edge Intersection} & 84.59 $\pm$ 0.61 & 77.06 $\pm$ 1.07 & 78.65 $\pm$ 0.94 & 74.86 $\pm$ 1.42 & 74.68 $\pm$ 0.83 & 73.45 $\pm$ 0.73\\
\bottomrule
\end{tabular}
\end{threeparttable}
}
\caption{Ablation study of the \methodname{} model over three datasets. We report the average and standard deviation over five runs, where Edge Intersection means taking intersection of edges from multiple dependency trees.}
\label{table:ablation}
\end{table}

\subsection{Results}

We first evaluate \methodname{} against previously established models, presenting the results in Table \ref{table:bestresults}. The \methodname{} model demonstrates superior performance on all datasets, particularly on the Laptop dataset, where it significantly outperforms existing models in both accuracy and Macro-F1 scores.

Performance comparisons of \methodname{} with baseline models are shown in Table \ref{table:results}. We observe that:

\textbf{Advantage of Syntax-based Models.} Dependency tree-based GAT and RGAT models consistently outperform the \textit{BERT-baseline}, underscoring the value of syntactic structure information in aspect-level sentiment classification.

\textbf{Benefits of Ensemble Approaches.} Ensemble models like \textit{Label-Ensemble}, \textit{Feature-Ensemble}, and \methodname{} surpass their single-tree counterparts. This highlights the advantage of leveraging multiple parses, reducing sensitivity to errors from any single parser.

\textbf{Superiority of \methodname{}.} Our \methodname{} model not only improves over single-tree models but also excels compared to other ensemble approaches, achieving the best performance without additional computational complexity.

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\begin{table}
\centering
\small
% \begin{adjustbox}{max width=0.475\textwidth}
\begin{tabular}{ c c c c }
\toprule
& Aspect Terms & Opinion Words & Coverage\\
\midrule
Laptop & \phantom{0}638 & 467 & 73.20\%\\
Restaurant & 1120 & 852 & 76.07\%\\
\bottomrule
\end{tabular}
% \end{adjustbox}
\caption{Statistics of the Opinion datasets. Aspect Terms denotes as the total number of aspect terms. Opinion Words represents total number of terms that have labeled opinion words. Coverage is the proportion of terms with labeled opinion words.}
\label{table:opinion_data}
\end{table}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\subsection{Model Analysis}

\textbf{Ablation Study.}
We conduct an ablation study to dissect the influence of model components, as shown in Table \ref{table:ablation}. Removing edge type or position embeddings leads to performance drops, indicating the importance of maintaining syntactic dependency information and reinforcing positional context in our model.

\textbf{Edge Union vs. Intersection.} We compare \methodname{}, which retains all edges from different trees, against an edge intersection approach that keeps only commonly shared edges. The latter, while reducing structural noise, underperforms \methodname{}, suggesting the benefit of the model's ability to navigate through the richer structural landscape offered by the ensemble graph.

\textbf{Aspect Robustness Assessment.} To evaluate the aspect robustness of our model, we utilize the ARTS datasets. Our model is tested against robustness variations in sentiment polarity, demonstrating better resilience compared to single dependency tree models. This indicates \methodname{}'s enhanced capability to handle complex sentiment variations in aspect-based sentiment analysis.

\section{Conclusion}

In this paper, we introduced \methodname{}, an innovative and straightforward graph-ensemble methodology, designed to amalgamate multiple dependency trees for enhanced aspect-level sentiment analysis. This technique, by integrating the collective edges from various parsers into a unified graph, empowers graph neural network models to withstand parsing inaccuracies effectively. This robustness is achieved without necessitating extra computational resources or an increase in model parameters, which is a significant advancement over existing methods.
The unique aspect of \methodname{} lies in its ability to preserve and utilize the inherent syntactic dependencies present in individual parse trees. By categorizing edges according to their original syntactic roles, our model not only retains but also enriches the contextual and structural understanding of the sentences under analysis. This nuanced approach to integrating syntactic information enables \methodname{} to deliver superior performance, outshining contemporary state-of-the-art models, single-parse approaches, and traditional ensemble methods. We demonstrated this across three benchmark datasets for aspect-level sentiment classification, each presenting its own set of challenges and linguistic nuances.
Our research presents a promising direction for future work in the field of sentiment analysis, particularly in leveraging syntactic structures more effectively. The success of \methodname{} in handling multiple dependency trees without additional computational burden opens up possibilities for further exploration into more complex and nuanced models that can better understand and interpret the subtleties of human language, especially in aspect-based sentiment contexts.

\bibliographystyle{unsrtnat}
\bibliography{references}

\end{document}